### v2

# Optimal Approximations for the Requirement Cut Problem on Sparse Graph Classes

Abstract
We study the Requirement Cut problem, a generalization of numerous classical graph partitioning problems including Multicut, Multiway Cut, $k$-Cut, and Steiner Multicut among others. Given a graph with edge costs, terminal groups $S_{1},\ldots,S_{g}$ and integer requirements $r_{1},\ldots,r_{g}$; the goal is to compute a minimum-cost edge cut that separates each group $S_{i}$ into at least $r_{i}$ connected components. Despite many efforts, the best known approximation for Requirement Cut yields a double-logarithmic $\mathcal{O}(\log(g)\cdot\log(n))$ approximation ratio as it relies on embedding general graphs into trees and solving the tree instance.
In this paper, we explore two largely unstudied structural parameters in order to obtain single-logarithmic approximation ratios: (1) the number of minimal Steiner trees in the instance, which in particular is upper-bounded by the number of spanning trees of the graphs multiplied by $g$, and (2) the 0ptof series-parallel graphs. Specifically, we show that if the number of minimal Steiner trees is polynomial in $n$, then a simple LP-rounding algorithm yields an $\mathcal{O}(\log n)$-approximation, and if the graph is series-parallel with a constant depth then a refined analysis of a known probabilistic embedding yields a $\mathcal{O}(0pt\cdot\log g)$-approximation on series-parallel graphs of bounded depth. Both results extend the known class of graphs that have a single-logarithmic approximation ratio.
keywords:
Requirement Cut, LP rounding, Randomized rounding, Spanning trees, Series-parallel graphs, Approximation algorithms, Graph partitioning
category:
[…]
1 Introduction
Graph partitioning problems are a central topic in combinatorial optimization, with applications such as from network design [14], VLSI Layout [10], image segmentation [23, 4], parallel computing [16] and even bioinformatics [29]. In this paper, we study the Requirement Cut problem, a generalization of several well-known cut problems. Given an undirected graph $G=(V,E)$ with edge costs, a collection of $g$ terminal groups $\mathcal{S}=(S_{1},\dots,S_{g})\subseteq V$, and a requirement $r_{i}>1$ for each group, the goal is to compute a minimum-cost edge cut $C\subseteq E$ that separates each terminal group $S_{i}$ into at least $r_{i}$ components in $G^{\prime}=(V,E\setminus C).$
The Requirement Cut problem is a natural extension of multiple classical problems that have been studied for decades. For each bullet point, next to the name of the problem, we will reference a paper that appears to be the one that introduced the problem to the community. Inside each bullet points, we will reference papers and results that have the current best general and optionally other more specific parametrized approximation ratio or run-times:
  • Multicut [27] : Given a set of terminal pairs $(s_{1},t_{1}),\dots,(s_{k},t_{k})$, the goal is to find a minimum-cost edge cut that separates each pair. This corresponds to the Requirement Cut problem when each terminal group consists of exactly two vertices and has a requirement of $r_{i}=2$. Garg et al. [21] gives the best known approximation for general graphs with a $\mathcal{O}(\log(k))$-approximation ratio via LP rounding (region growing). Friedrich et al. give a constant approximation for series-parallel graphs [19] and a more general [20] $\mathcal{O}(\log(tw))$ for graphs of treewidth $tw$.
  • Multiway Cut [12]: Given a set of terminals $S=\{s_{1},\dots,s_{k}\}$, the objective is to find a minimum-cost edge cut that separates all terminals into distinct components. This is a special case of Requirement Cut where there is only one group and $r=k$. Chopra et al. [9] studied the linear program of this problem in depth before Calinescu et al. [11] improved the best approximation from $2-\frac{2}{k}$ to $1.5-\frac{1}{k}$. This problem also has seen a relatively new variant in the past few years called the Norm Multiway Cut[7, 6].
  • Steiner-Multicut [31]: Given sets of terminals $\mathcal{S}=(S_{1},\dots,S_{k})$ each of size $(t_{1},\dots,t_{k})$, the goal is to find a minimum-cost edge cut that separates each set in at least 2 components. This corresponds to the Requirement Cut problem when each requirement is 2. This problem has particularly many applications has even been studied in a parametrized complexity setting [5]. Its best current approximation comes directly from the Requirement Cut and comes with a $\mathcal{O}(\log(k)\log(n))$ ratio. Moreover, it is a direct generalization of a less studied variant when $k=1$, the Steiner-Cut problem [28].
  • Multi-Multiway Cut [2]: A generalization of Multiway Cut where multiple terminal groups must be separated into distinct connected components. This corresponds to the Requirement Cut problem when each terminal group has a requirement equal to its size. The best known approximation comes from the paper that introduced the problem [2] and is $\log(2k)$. Slightly later two more papers have studied the problem under different angles, [35] by adding label edges and [13] studying the problem on bounded branch-width graphs.
  • k-Cut [22] : The problem of partitioning a graph into at least $k$ connected components using a minimum-cost edge cut. This can be viewed as a Requirement Cut instance where there is only $S_{1}=V$ as a terminal set and $r_{1}=k$. After the introduction of the problem that studied its parametrized complexity, Saran and Vazirani[34] found the current best approximation ratio $2-\frac{2}{k}$. More recently, Gupta et al. [24], gave tight bounds for the run-time of computing an optimal solution.
  • Steiner k-Cut [8]: This is a generalization of the k-Cut problem, where $S_{1}\subseteq V,|S_{1}|\geq k$ and the objective is to find a minimum cost set of edges whose removal results in at least $r_{1}=k$ disconnected components, each containing a terminal. When $k=n$ this is exactly the k-Cut problem. The problem has been established and approximated best by Chekuri et al. [8] to a $2-\frac{2}{k}$ ratio. It is worth noting that their work has established the Linear Program formulation used to solve the Requirement Cut Problem introduced later.
Refer to caption[21][20][11][34][25][2][8][25](Section 3)(Section 5)
Figure 1: Hierarchy of graph cut problems captured by the Requirement Cut framework putting our results in perspective. Boxes contain parameters, and best known approximation results for general graphs (and for more particular cases when specified). Arrows denote special cases of the problems (i.e., the source is a particular case of the target).
These problems have been studied mostly through the lens of approximation algorithms. Some of them have received particular attention on specific classes of graphs with parameters such as treewidth, branchwidth or pathwidth.
The hardness of the Requirement Cut can be shown through any of these problems since they are all hard and are particular cases of it. A particular reduction that helps us with bounding the approximability on trees is a reduction from Set Cover taken from [32]. Consider a reduction from Set Cover to the Requirement Cut problem on trees. We construct a star graph where each leaf node corresponds to a set in the Set Cover instance, and the center node of the star connects to each of these leaves via an edge of cost $1$. For each element $j$ in the universe, we define a group consisting of the root and all leaf nodes whose corresponding sets contain element $j$. We assign a requirement of $2$ to each group. In this construction, selecting an edge in the tree (i.e., choosing a set in the Set Cover instance) ensures that the corresponding group is cut (i.e., its requirement is satisfied) if and only if that set covers one of the group’s elements. Hence, there is a one-to-one correspondence between feasible solutions of the Requirement Cut instance and valid set covers, and this correspondence preserves the total cost. Feige et al. [18] showed that Set Cover can not be approximated in polynomial time to a $\Omega(\log(n))$ ratio where $n$ is the size of the universe to cover. It follows that Requirement Cut on trees is hard to $\Omega(\log(g))$- approximate where $g$ is the number of groups.
Now that we have seen that the best approximation possible on trees for our problem is $\log(g)$, let us put the results of our paper in context. Note that, the reduction from Set Cover to Requirement Cut on trees works even when all the requirements are 2 (i.e. the Requirement Cut instance is a Steiner Multicut instance). In 1997, Klein, Plotkin, Rao and Tardos [31] presented an algorithm that yields $\mathcal{O}(\log^{3}(gt))$-approximation with $t=\max(t_{i})$ (the sizes of the sets) for the Steiner Multicut problem. This “triple log” approximation ratio remained the best for this problem until 2010. In 2010, the Requirement Cut problem was introduced by Nagarajan and Ravi [32] who proved a $\mathcal{O}(\log(n)\cdot\log(gR))$-approximation with $R=\max(r_{i})$. This approximation became, at the same time, the best approximation ratio for the Steiner Multicut problem. Their bound was later improved by the same set of authors along with Gupta [25] that got the approximation ratio down to $\mathcal{O}(\log(n)\cdot\log(g))$, effectively removing all dependency on the requirements. Both these approaches, that manage to reduce the power of “logs” in the approximation ratio from triple to double, involve very similar algorithmic steps:
  1. 1.
    Finding a good approximation algorithm for trees, $\log(gR)$ for [32] and $\log(g)$(tight) for [25].
  2. 2.
    Using Fakcharoenphol, Rao and Talwar (FRT) [17], they embed the graph into a tree and pay a $\log(n)$ factor for it.
  3. 3.
    They use the the first approximation on the resulting tree and get their $\mathcal{O}(\log(gR)\cdot\log(n))$ or $\mathcal{O}(\log(g)\cdot\log(n))$-approximation.
It is also worth noting that Gupta et al. [25] technically has a slightly better $\mathcal{O}(\log(k)\cdot\log(g))$-approximation with $k$ being the total number of terminals. They obtain this improvement through something that they call a slight generalization of FRT [17] in which they embed into a tree only the terminal vertices and hence only pay $\log(k)$ and not $\log(n)$. But this remains secondary in their paper since their main contribution was closing the gap between the $\log(gR)$ approximation on trees and the known lower bound of $\log(g)$ making it tight.
For the past few years, one goal of the community has been to beat those “double log” approximations with “single log” ones. Since we know that embedding into trees has a $\log(n)$ cost, (surprisingly) even for treewidth 2 graphs [26], and that a Requirement Cut problem instance on trees can not be approximated better than with a $\log(g)$ factor [18], we know that continuing this path of using tree embeddings will not result into a “single log” approximation. In this paper, we try something different over two different routes that we present in the next section.
1.1 Our Contributions
We address the following question: For which classes of graphs does the Requirement Cut problem admit optimal (single-log) approximation algorithms? To this end, we identify two structural parameters—(1) the number of spanning or minimal Steiner trees of the input graph and (2) the depth of series-parallel graphs — we show that both these relatively unstudied parameters offer better approximation for the Requirement Cut problem.
First, we show that for series-parallel graphs of constant 0pt, a refined analysis of Emek and Peleg’s [15] existing embeddings method yields an $\mathcal{O}(\log(g))$-approximation, extending the known $\mathcal{O}(log(g))$-approximations for trees to a broader class of graphs.
Second, we demonstrate that graphs with a polynomial number of spanning trees (or, more generally, number of minimal Steiner trees) admit an $\mathcal{O}(\log(n))$-approximation via a randomized LP-rounding scheme similar to the one given by Nagarajan and Ravi [32].
Our results broaden the landscape of instances that can be approximated with a “single logarithmic” factor and shed light on two unstudied parameters until now, $depth$ of series-parallel graphs and the order of magnitude of the number of spanning trees in the graph in the context of Requirement Cut.
  • Instances with a Polynomial Number of Steiner Trees: A key parameter in our analysis of our Algorithm 1 of Section 4 depends on the instance. We will refer to it as $\sigma(G,\mathcal{S})$ for a given Requirement Cut instance. This function is equal to the number of distinct Minimal Steiner Trees (Steiner Trees for which all the leaves are terminal). This quantity is hard to compute [36] and so we will almost always upper-bound it by $\tau(G)\cdot g$ with $\tau(G)$ being the number of distinct spanning trees of the graph which is computable in polynomial time using Kirchhoff’s Matrix Tree Theorem initially stated in [30]. We will show that if $\tau(G)=\mathcal{O}(n^{c_{1}})$ then necessarily $\sigma(G,\mathcal{S})=\mathcal{O}(n^{c_{2}})$ when $c_{1}$ and $c_{2}$ are constants.
  • Bounded-Depth Series-Parallel Graphs: Series-parallel graphs are well known for their recursive structure. They are often studied as a generalization of trees since they are, for example, representative of treewidth 2 graphs (trees have treewidth 1) and also are exactly $K_{4}$-minor-free graphs (trees are $K_{3}$-minor-free). In our paper, we introduce their depth as a key parameter. This parameter has never been studied to the best of our knowledge and we show a $\mathcal{O}(depth\cdot\log(g))$-approximation for the Requirement Cut problem on series-parallel graphs. Extending the class of graphs that have a $\mathcal{O}(\log(g))$-approximation to the series-parallel graph with any constant $depth$. This class of graph can additionally be seen as a generalization of melonic graphs which are used in other fields such as physics.
In Section 3 we will give several classes of graphs and instances for which our results apply. Afterwards in Section 4 we give, prove and analyze our algorithm regarding graphs of polynomial number of Steiner trees. This comes with a proof about the Linear Program formulation of the Requirement Cut problem that is of its own interest (Lemma 4.1). Finally in Section 5, we refine Emek and Peleg’s [15] analysis to adapt it to the depth of series-parallel graphs. This approach provides an alternative to the known $\log(n)$ embeddings for series-parallel graphs.
2 Preliminaries
We begin by defining the key concepts used throughout the paper.
Definition 2.1 (Cut).
Let $G=(V,E)$ be an undirected graph. A cut in $G$ is a set of edges $C\subseteq E$ such that removing the edges in $C$ disconnects some specified structure in the graph. We say that a cut separates a set of vertices $S\subseteq V$ if the vertices in $S$ lie in at least two distinct connected components of $G\setminus C$.
Definition 2.2 (Requirement Cut Instance).
A Requirement Cut instance is defined by an undirected graph $G=(V,E)$, a cost function $c:E\rightarrow\mathbb{R}_{\geq 0}$, a collection of terminal sets $\mathcal{S}=\{S_{1},\ldots,S_{g}\}$ with $S_{i}\subseteq V$, and integer requirements $r_{1},\ldots,r_{g}\geq 2$. The goal is to find a minimum-cost edge set $C\subseteq E$ such that in the graph $G^{\prime}=(V,E\setminus C)$, each terminal set $S_{i}$ is partitioned into at least $r_{i}$ connected components.
For the two next definitions, let $G=(V,E)$ and $\mathcal{S}=\{S_{1},\ldots,S_{g}\}\subseteq V$ a set of terminal sets.
Definition 2.3 (Steiner Tree).
A Steiner tree for a terminal set $S$ is a minimally connected subgraph $T\subseteq G$ that spans all vertices in $S$. $T$ may include additional non-terminal vertices, called Steiner vertices. An edge connecting two terminals (resp. Steiner) vertices may be called terminal edge (resp. Steiner edge).
Definition 2.4 (Minimal Steiner Tree).
A Steiner tree is said to be minimal if all its leaves are terminal vertices. The number of distinct minimal Steiner trees in a graph $G=(V,E)$ with terminal sets $\mathcal{S}=\{S_{1},\ldots,S_{g}\}$, will be noted as $\sigma(G,\mathcal{S})$.
With this definition of Minimal Steiner tree, let us prove the following claim:
Claim 1.
In a graph $G=(V,E)$ with $\mathcal{S}=\{S_{1},\ldots,S_{g}\}$ terminal sets, we have : $\sigma(G,\mathcal{S})\leq g\cdot\tau(G)$.
Proof 2.5.
Let $T$ be a minimal Steiner tree of set $S_{k}$ for some $1\leq k\leq g$. By adding edges and vertices to $T$ until it covers $V(G)$ one can create at least one spanning tree per minimal Steiner tree. Also two distinct minimal Steiner trees (even of the same terminal set) can not extend to the same spanning tree as they differ in at least one edge.
Definition 2.6 (Series-Parallel Graph).
The class of series-parallel graphs is recursively defined as follows.
  • The graph consisting of a single edge $(x,y)$ is series-parallel with terminals $x$ and $y$.
  • Let $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ be two series-parallel graphs with terminals $x_{1},y_{1}\in V_{1}$ and $x_{2},y_{2}\in V_{2}$. Then:
    • The series composition of $G_{1}$ and $G_{2}$ identifies $y_{1}$ with $x_{2}$ (creating a new vertex $z$), and results in a new graph with terminals $x=x_{1}$ and $y=y_{2}$.
    • The parallel composition identifies $x_{1}$ with $x_{2}$ (creating $x$), and $y_{1}$ with $y_{2}$ (creating $y$), resulting in a new graph with terminals $x$ and $y$.
Definition 2.7 (Composition Trace).
Let $G$ be a series-parallel graph constructed via a sequence of series and parallel compositions. The composition trace of $G$ is a directed tree $T$ whose:
  • leaves correspond to the individual edges of $G$,
  • internal nodes represent subgraphs formed by (series or parallel) composition of their children
  • root corresponds to $G$ itself.
We group consecutive series or parallel compositions into single nodes, so that every path from a leaf to the root alternates between series and parallel composition steps. A node in $T$ may have arbitrarily many children.
Definition 2.8 (Depth of a Series-Parallel Graph).
Let $G$ be a series-parallel graph and let $T$ be its composition trace. The depth of $G$, is defined as the depth of the tree $T$, that is, if $l$ is the length of the longest path from a leaf (corresponding to a single edge) to the root (corresponding to $G$ itself): 0pt$=l-1$ Equivalently, it is the number of composition steps along the longest sequence of nested series and parallel compositions used to construct $G$.
For example, a single edge would be a series-parallel graph of depth 0, a path or a many parallel edges would have depth 1 and a composition of parallel paths would have depth 2.
3 Graphs with Polynomial Number of Steiner Trees
In this section, we will see for which graphs our Algorithm 1, introduced later in Section 4, is more efficient than the state of the art. First, we will comment on the link between the number of spanning trees and the number of Steiner trees, and then we will also show that the number of spanning trees is not really limiting our algorithm and that there are other graphs where our algorithm provides an efficient approximation as well.
[…]
Figure 2: Overview of graph classes and instances for which the Requirement Cut problem admits an improved (single log) approximation in our paper.
3.1 Graphs with Polynomial number of Spanning Trees
A key parameter in our analysis is the number of spanning trees in a graph, which we denote as $\tau(G)$. By Kirchhoff’s Matrix Tree Theorem [30], $\tau(G)$ can be computed as the determinant of any cofactor of the graph Laplacian matrix of the graph. Equivalently, if $\lambda_{2},\lambda_{3},\dots,\lambda_{n}$ are the nonzero eigenvalues of $L$, then $\tau(G)=\frac{1}{n}\prod_{i=2}^{n}\lambda_{i}$. In our setting, we focus on graphs where $\tau(G)$ is polynomial in $n$, i.e., $\tau(G)=\mathcal{O}(n^{c})$ for some constant $c$.
The number of spanning trees multiplied by $g$ (the number of terminal sets of our Requirement Cut problem instance) is a natural upper bound to the number of distinct minimal Steiner trees in that same instance. So naturally, graphs with a polynomial number of spanning trees satisfy our condition. Let us now see three examples of graph classes that satisfy our spanning tree condition.
3.1.1 Graphs with Small Feedback Edge Sets
One prominent category of graphs with polynomially many spanning trees is the graphs that have a small, or bounded, feedback edge set. The feedback edge set is defined as the minimum number of edges whose removal results in an acyclic graph. Note that the feedback edge set number of a connected n-vertex and m-edge graph G is always $m-n+1$ and a minimum feedback edge set can be determined by the computation of a spanning tree in $\mathcal{O}(n+m)$ time via depth-first search. Also note that this class of graphs contains graphs with bounded number of cycles.
Back to our case, if the size of a graph’s feedback edge set is bounded by a constant $k$, then the number of spanning trees is $\binom{n-1+k}{k}=\mathcal{O}(n^{k})$, and hence polynomially bounded in the size of the graph.
3.1.2 Graphs with Short Cycles
Graphs with short cycles, meaning each cycle contains up to $k$ vertices and the number of cycles is of order $\mathcal{O}(\log(n))$, have a polynomial number of spanning trees. In such a graph, a spanning tree is obtained by removing $1$ edge for each cycle making the number of spanning trees in the order of $\mathcal{O}(k^{\log(n)})=\mathcal{O}(n^{\log(k)})$ which is polynomial for fixed k.
[…]
Figure 3: This graph has $\mathcal{O}(\log(n))$ cycles of length 3. each spanning tree chooses 1 edge per cycle to remove, hence yielding a number of spanning trees in $\mathcal{O}(3^{\log(n)})$ which is polynomial $\mathcal{O}(n^{\log(3)})$.
3.2 Instances of Polynomial number of Steiner Trees
The Steiner trees in which we are interested for our algorithm are minimal Steiner trees, as given by Definition 2.4. We denote the number of such trees for a given instance by $\sigma(G,\mathcal{S})$. This value depends on the terminal sets of the instance as well as on the graph, and is hence more specific than its upper bound $\tau(G)\cdot g$. Unfortunately it is hard to compute. Still, it can provide better approximation for some type of instances where the $\tau(G)$ would yield a value that is super-polynomial in $n$.
Those instances are the instances where the whole graph may have a super-polynomial number of spanning trees but the number of Steiner trees is limited. This can occur for example if the terminal sets are each fully located in subgraphs that have a polynomial number of spanning trees when cut from the rest of the graph and that edge-cut to the rest of the graph is of size 1.
[…]
Figure 4: Instance where terminal groups reside in minimally connected subgraphs. Here, it would be enough that $\tau(H_{i})=poly(|H_{i}|),i\in\{1,2,3\}$ to have a $\log(n)$ approximation for the instance.
4 Requirement Cut Algorithm
4.1 Linear Program
We work with essentially the same $LPs$ as [32] and [25]. However, in Lemma 4.1, we will provide additional insights about these linear programs, compared to previous works. This is central to our analysis as it allows us to guarantee that our rounded solution is indeed a valid solution to the Requirement Cut problem. Both these $LPs$ work on slightly modified graphs. The first, IP-Steiner, calls for a complete graph, so we add 0-cost edges to the graph until it is a clique, we call this graph completed $G_{V}$. The second, Relaxed LP, calls for extra edges only between terminals of the same group. We shall call this modified graph, where $S_{i}$ is a clique with the addition of 0-cost edges, $G_{S_{i}}=(V,E_{S_{i}})$. This second modification is slightly different to the one used by [32] and [25] who used the induced graph $G[S_{i}]$ for the first and the complete graph on $S_{i}$ ($G=(S_{i},\binom{S_{i}}{2})$for the other. Our relaxation is equivalent when it comes to the purpose of this paper.
[…]
s.t. $\sum_{e\in\mathcal{T}_{i}}d_{e}\geq r_{i}-1\quad$ $\forall\mathcal{T}_{i}:S_{i}\text{-Steiner tree in }G_{V},\quad\forall i=1\dots g$
$d_{e}\in\{0,1\}\quad$ $\forall e\in E$
(Relaxed $(\text{Relaxed LP})\quad\min$ $\sum_{e\in E}c_{e}d_{e}$
s.t. $\sum_{e\in\mathcal{T}_{i}}d_{e}\geq r_{i}-1\quad\forall\mathcal{T}_{i}:\text{spanning tree in }G_{S_{i}},\forall i=1\dots g$
$d_{\{u,w\}}\leq d_{\{u,v\}}+d_{\{v,w\}}\quad\forall u,v,w\in V$
$0\leq d_{\{u,v\}}\leq 1\quad\forall u,v\in V$
IP-Steiner is an exact formulation of the Requirement Cut problem. However, we can notice that it is not clear why with only the constraint on the number of cut edges per Steiner tree, the solution would be a solution to the Requirement Cut problem. In Lemma 4.1 we will precisely show the role of the completeness of the graph in IP-Steiner. In fact, without the completeness of the graph in IP-Steiner (or of the terminal sets in Relaxed LP), Steiner trees could have a long $d$ length without satisfying the requirement. For example, take a path that starts and ends with terminal vertices and arbitrarily many Steiner vertices in the middle. This graph could be cut $r-1$ times satisfying the constraint on Steiner trees but having its terminal lying only in two components and not $r$.
The algorithm we present in this section is a randomized rounding of the Relaxed LP solution which can be solved by an ellipsoid method, using an MST algorithm as an oracle. What it ensures is that each and every Steiner tree is cut at least $r-1$ times. This alone does not suffice for proving that it returns a valid solution. Actually in [32] Nagarajan and Ravi use a preprocessing of their graph allowing them to have a relation between number of edges removed from the graph and number of terminal components created (Claim 2 in their paper). This same claim is used in Gupta, Nagarajan and Ravi [25] (Claim 3 in that paper) and it can be used simply because they are working with trees. In our case, we need something slightly stronger and more general since we are working with a more general class of graphs. Hence, before presenting our algorithm we will show an equivalence (under the condition of 0-cost edges added of our $LP$) between the two statements “All terminals of a set $S_{i}$ are in $r_{i}$ components” and “All spanning trees in $G_{S_{i}}$ are cut at least $r_{i}-1$ times.” with $G_{S_{i}}$ being the graph $G$ with extra 0-cost edges making the vertices of $S_{i}$ a clique. Also, just like the two papers mentioned, let $d^{*}$ be the solution to Relaxed LP, we work with $d_{\{u,v\}}=\min(2\cdot d^{*}_{\{u,v\}},1)$ which is also a metric since $d^{*}$ is one.
Claim 2 ([32]).
For any group $S_{i},i\in(1,\dots,g)$, the minimum Steiner tree on $S_{i}$ under metric $d$ has length at least $r_{i}-1$.
Lemma 4.1.
Let $G=(V,E)$ a graph, $\mathcal{S}=\{S_{1},\ldots,S_{g}\}$ with $S_{i}\subseteq V$ a collection of terminal sets, and integer requirements $r_{1},\ldots,r_{g}\geq 2$. Let each $G_{S_{i}}$ be the graph where $S_{i}$ is connected by 0-cost edges making it a clique. Let $d:E\rightarrow\{0,1\}$ a metric such that for every $S_{i}$-Steiner tree $T$, $d(T)\geq r_{i}-1$. Then it is equivalent to say that all spanning trees over $G_{S_{i}}$ are cut at least $r_{i}-1$ times and to say that the vertices of $S_{i}$ lie in $r_{i}$ components.
Proof 4.2.
For the purpose of this proof we drop the indexes by considering the graph $G_{S}=(V,E_{S})$ with $S$ being a terminal set with requirement $r$. It is enough to show the property for one terminal set since more terminal sets only cut the graph in more components by cutting more edges. We call an edge $e$ with $d(e)=1$ a cut-edge. We call a component containing a terminal vertex, a terminal component. A non-terminal component is a component that does not contain a terminal vertex.
Let $T_{0}$ be an $S$-Steiner tree in $G$. $d(T_{0})\geq r-1$ and $S$ in less than $r$ components of $T_{0}=(k_{1},\dots,k_{\chi})$ with $\chi\geq r$. Let $k^{*}\in T_{0}$ be a non-terminal component. We know that $k^{*}$ is not a leaf component since $T_{0}$ is a minimal Steiner tree and hence it has at least 2 cut-edges incident to it, we note $c(k^{*})\geq 2$ this number. Also, let $t_{1},t_{2}\subseteq T_{0}\setminus k^{*}$ be two terminal subtrees adjacent to $k^{*}$. Let $x\in t_{1}$ and $y\in t_{2}$ two terminal vertices and $e^{*}$ the 0-cost edge in $G_{S}$ connecting them. Now we have $T_{1}$ another $S$-Steiner tree with $T_{1}=\{e^{*}\}\cup T_{0}\setminus k^{*}$. By our construction $d(T_{1})\geq d(T_{0})-c(k^{*})+1$ the $+1$ being $e^{*}$, by property of $d$, $d(T_{1})\geq r-1$. Now we have 2 case:
  1. 1.
    $d(T_{0})-c(k^{*})+1<r-1$: Then $T_{1}$ need to have a cut in one of the terminal subtrees effectively increasing its terminal components (along with $T_{0}$’s) by $1$ because it enters $t_{1}$ and $t_{2}$ at a terminal. We can then iterate over other non terminal components in $T_{0}\cup T_{1}$.
  2. 2.
    $d(T_{0})-c(k^{*})+1\geq r-1$: Then $T_{1}$ does not need anymore cut edges. In this case we have two subcases:
    1. (a)
      In $T_{1}$, $S$ lies in $r$ components: This contradicts the assumption since this would mean that $T_{0}$ lied in $r$ components.
    2. (b)
      In $T_{1}$, $S$ lies in less than $r$ components and so we have the same procedure creating $T_{2}$ that shortcuts one of the non-terminal components of $T_{1}$.
For case $(1)$ we can see that it increases the number of terminal components until we have created enough in some $T_{i}$ down the exploration.
For case $2b$ we have eliminated $k^{*}$ from $T_{1}$ and we can now attack and shortcut another non terminal component. This will necessarily converge since at each iteration we reduce by at least one the $d(T_{i})$.
This achieves the proof of the lemma.
Now that we proved Lemma 4.1, we can show our Algorithm that will ensure that each Steiner tree is cut with high probability enough times gives us a solution to the Requirement Cut problem with the same probability.
4.2 Algorithm
In this section we will see how our algorithm can give us a $\mathcal{O}(\log(\sigma(G,\mathcal{S})))$-approximation for the Requirement Cut problem where $\sigma(G,\mathcal{S})$ is the number of distinct minimal Steiner trees in the graph with respect to the given terminal sets. Computing $\sigma(G,\mathcal{S})$ is #$\texttt{\#}P$-Complete is since it is a more general case of counting paths [36]. But, thankfully, we can upper-bound it by $\tau(G)\cdot g$ (1) which is easy to compute through Kirchhoff’s Theorem [30]. When clear from the context, we shall denote the number $\sigma(G,\mathcal{S})$ simply by $\sigma$.
Algorithm 1 Randomized Rounding for the Requirement Cut problem
1:Input: Instance of the Requirement Cut problem on graph $G=(V,E)$
2:Solve the LP relaxation to obtain fractional values $\{d(e)\in[0,1]\mid e\in E\}$
3:Initialize cut set $C\leftarrow\emptyset$
4:Choose $\alpha=\frac{1}{c\cdot\log(\sigma)}$ with $c$ large enough (see Section 4.2)
5:for each edge $e\in E$ do
6:     Draw $x_{e}\sim\text{Uniform}(0,\alpha)$
7:     if $x_{e}\leq d(e)$ then
8:         Add $e$ to $C$
9:     end if
10:end for
11:Output: Cut set $C$
The rest of this section of the paper will be dedicated to proving this theorem:
Theorem 4.3.
Let $I$ be a Requirement Cut Instance (Definition 2.2) with $\sigma$ being the number of minimal Steiner trees in $I$. Then, there is an LP rounding algorithm for the Requirement Cut problem that returns a solution of cost $\mathcal{O}(\log(\sigma))\cdot LP_{opt}$ with probability $1-\frac{1}{poly(\sigma)}$, where $LP_{opt}$ is the value returned by Relaxed LP.
In other words the theorem provides $\mathcal{O}(\log(n))$-factor approximation for all graphs described in (Section 3) and others. It also matches the $\mathcal{O}(\log(g))$-approximation on trees given by [25]. We believe that our analysis is also simpler, since it only involves one of the two phases described in [25] (i.e., the second one).
Lemma 4.4.
The expected cost of $C$, the set of edges returned by Algorithm 1 is $\mathds{E}(c(C))\leq\frac{1}{\alpha}\sum_{e\in E}c_{e}d_{e}=\mathcal{O}(\log(\sigma))\cdot LP_{opt}$.
Proof 4.5.
The probability of a given edge $e$ being in the cut $C$ is $\mathbb{P}(e\in C)=\min\left\{1,\frac{d(e)}{\alpha}\right\}$. Thus, the expected cost of $C$ is $\mathds{E}(c(C))\leq\frac{1}{\alpha}\sum_{e\in E}c_{e}d_{e}=\mathcal{O}(\log(\sigma))\cdot LP_{opt}$ by linearity of expectations.
Lemma 4.6.
After the execution of Algorithm 1 the probability that a group $S_{i}$ lies in less than $r_{i}$ is $\frac{1}{\sigma^{\frac{c}{2}-1}}$ with $c\geq 4$ a constant chosen arbitrarily.
We will call a terminal set $i$ that lies in at least $r_{i}$ components satisfied.
Proof 4.7.
Let $\alpha=\frac{1}{c\cdot\log(\sigma)}$ with $c$ a constant chosen large enough (Section 4.2) and define $x_{e}\sim\text{Uniform}(0,\alpha)$ picked independently for each $e\in E$. Let $d:E\to[0,1]$ be the solution returned by Relaxed LP, and let $\mathcal{T}_{i}$ be the set of all Steiner trees of group $g_{i}$ in $G$. By LP feasibility, we have:
$\sum_{e\in t}d(e)\geq r_{i}-1,\quad\forall t\in\mathcal{T}_{i}$ (1)
We define an indicator variable $Z_{e}$ for each edge $e\in E$ as follows:
$Z_{e}=\begin{cases}0,&\text{if }x_{e}>d_{e},\\ 1,&\text{otherwise}.\end{cases}$
Next, define a random variable $Y_{t}$ for any Steiner tree $t\in\mathcal{T}_{i}$:
$Y_{t}=\sum_{e\in t}Z_{e}.$
Claim 3.
For all $t\in\mathcal{T}_{i}$, we have $\mathds{E}[Y_{t}]\geq r_{i}-1$. In other words, the expected number of edges cut in each Steiner tree is at least $r_{i}-1$.
Proof 4.8.
We rely on Equation 1 and the probability that any edge $e$ is cut is given by
$\mathbb{P}(Z_{e}=1)=\min\left\{1,\frac{d(e)}{\alpha}\right\}$
Taking expectations, we obtain:
$\mathds{E}[Z_{e}]\geq d(e).$
Summing over all edges in $t$, it follows that
$\mathds{E}[Y_{t}]=\sum_{e\in t}\mathds{E}[Z_{e}]\geq\sum_{e\in t}d(e)\geq r_{i}-1.$
This completes the proof of 3.
Now that we have established that each Steiner tree is expected to be cut at least $r_{i}-1$ times, we analyze the probability that every $S_{i}$-Steiner tree is cut $r_{i}-1$ times. Proving that our algorithm returns a solution using Lemma 4.1.
For this analysis, we shall first do a few adjustments into our instance:
  1. 1.
    Separate into two sets the edges of our cut $C=C_{1}\uplus C_{2}$. On the one hand, $C_{1}$ is the set of edges such that $d(e)\geq\alpha\iff\mathbb{P}(Z_{e}=1)=1$. $C_{2}$, on the other hand, is the set of edges for which $d(e)<\alpha$, meaning they had a probability $\mathbb{P}(Z_{e}=1)=\frac{d(e)}{\alpha}$ of being picked in the cut.
  2. 2.
    For each set $S_{i}$, after the removal of the edges of $C_{1}$, some of its requirement might be satisfied. So we call $r_{i}^{\prime}$ the residual requirement after the removal of the edges in $C_{1}$. We have that $r_{i}^{\prime}\leq r_{i}$.
The goal of the next steps of the analysis will be to show that the edges in $C_{2}$ cover the residual requirements $r_{i}^{\prime}$ left after the removal of the edges in $C_{1}$ with high probability.
Since we are now only looking at the new graph $G=(V,E\setminus C_{1})$, let us update some of our bounds from before.
$\forall e\in E\setminus C_{1},\mathbb{P}(Z_{e}=1)=\mathds{E}[Z_{e}]=\frac{d(e)}{\alpha}$
By linearity of expectations we get :
$\forall i,\forall t\in\mathcal{T}_{i},\mathds{E}[Y_{t}]=\sum_{e\in t}\mathds{E}[Z_{e}]\geq\sum_{e\in t}\frac{d(e)}{\alpha}\geq\frac{r^{\prime}_{i}-1}{\alpha}$
Now let us bound the probability of one Steiner tree not being satisfied using a Chernoff bound. For convenience, we drop the indexes on the variables since the calculations hold for every Steiner tree $t$ in $G=(V,E\setminus C_{1})$.
We use the following Chernoff bound (the details for some of the steps are available in Section A.1):
$\ \mathbb{P}\left(Y_{t}<(1-\delta)\mathds{E}[Y_{t}]\right)\leq e^{-\frac{\delta^{2}}{2}\mathds{E}[Y_{t}]},\quad\text{for }0<\delta<1.$
Substituting $\delta=1-\frac{r^{\prime}-2}{\mathds{E}[Y_{t}]}$, we derive:
$\mathbb{P}(Y_{t}\leq r^{\prime}-2)\leq e^{-\frac{(1-\frac{r^{\prime}-2}{\mathds{E}[Y_{t}]})^{2}}{2}\mathds{E}[Y_{t}]}.$
After developping the exponent, we get:
$\exp({-\frac{(\mathds{E}[Y_{t}]-r^{\prime}-2)^{2}}{2\mathds{E}[Y_{t}]}})$
Using the lower bound on $\mathds{E}[Y_{t}]$ to replace it and simplifying:
$\exp\left(-\frac{(r^{\prime}-1-\alpha(r^{\prime}+2))^{2}}{2\cdot\frac{r^{\prime}-1}{\alpha}\cdot\alpha^{2}}\right)=\exp\left(-\frac{(r^{\prime}-1-\alpha(r^{\prime}+2))^{2}}{2\alpha(r^{\prime}-1)}\right)$
Finally, we know that $r^{\prime}\geq 2$.
$\exp\left(-\frac{(r^{\prime}-1-\alpha(r^{\prime}+2))^{2}}{2\alpha(r^{\prime}-1)}\right)\leq\exp\left(-\frac{(1-4\alpha)^{2}}{2\alpha}\right)$
$\leq\exp\left(-\frac{16\alpha^{2}-8\alpha+1}{2\alpha}\right)$
Substitute $\alpha$ for $\frac{1}{c\cdot\log(\sigma)}$:
$\exp\left(-\frac{\frac{16}{(c\cdot\log(\sigma))^{2}}-\frac{8}{c\cdot\log(\sigma)}+1}{\frac{2}{c\cdot\log(\sigma)}}\right)=\exp\left(-\frac{\frac{16}{(c\cdot\log(\sigma))}-8+c\cdot\log(\sigma)}{2}\right)$
Here we simply distribute the division by two and then remove a negative term.
$=\exp\left(4-\frac{8}{c\cdot\log(\sigma)}-\frac{c\cdot\log(\sigma)}{2}\right)\leq\exp\left(4-\frac{c\cdot\log(\sigma)}{2}\right)$
$=\frac{\exp(4)}{\sigma^{\frac{c}{2}}}$
Since $\sigma$ is $poly(n)$ this shows that for each $t\in\mathcal{T}_{i}$ the probability of not being satisfied is inverse polynomial in $n$.
Then the probability of $C$, the output of the algorithm, not being a solution is at most:
$\sum_{S_{i}\in\mathcal{S}}\sum_{e\in\mathcal{T}_{i}}\frac{\exp(4)}{\sigma^{\frac{c}{2}}}=\frac{\exp(4)\cdot\sigma}{\sigma^{\frac{c}{2}}}=\frac{\exp(4)}{\sigma^{\frac{c}{2}-1}}$
Under the condition that $\sigma=poly(n)$, this means that $C{}$ is a solution with high probability i.e. $1-\frac{1}{poly(n)}$.
Choosing alpha
Let us now discuss the details of fixing $\alpha$ or rather $c$. For some part of our analysis we need $c\geq 4$ assuming $\log(n)\geq 1$. Technically speaking, one could have a $\log(\sigma)$-approximation even for instances where $\sigma$ is not $poly(n)$. But this would not lead to a “single-log” approximation, as the factor is only $\mathcal{O}(\log(n))$ when $\sigma=poly(n)$ or smaller.
5 Depth of Series Parallel Graphs
In this section, we refine our approximation bounds for the Requirement Cut problem on series-parallel graphs by analyzing differently the algorithm given by Emek and Peleg[15]. This different analysis coupled with the $\mathcal{O}(\log(g))$-approximation for trees [25] yields a $\mathcal{O}(m\cdot\log(g))$-approximation for series-parallel graphs of depth $m$. Series-parallel graphs in general cannot be embedded into trees with less than $\log(n)$ expected distorsion [26]. This means that general series-parallel graphs currently have the same approximation ratio for our problem as general graphs, i.e., $\mathcal{O}(\log(g)\cdot\log(n))$. In fact, the infinite class of series-parallel graphs that Gupta et al. [26] introduce as having distortion no less than $\log(n)$ when embedded into trees have a 0ptof $\mathcal{O}(\log(n))$. Meaning that our result provides a credible alternative also in some cases when the 0ptis smaller than $\log(n)$ and not only when it is a given constant.
Refer to captiondepth $=3$depth $=5$depth $=7$
Figure 5: Series-parallel graphs of 0pt3, 5, and 7. These depths correspond to the number of nested compositions in the graph’s construction and form a subclass of graphs with bounded tree-depth, enabling single-log approximations for Requirement Cut. Note that the number of paths in those graphs can be exponential and that each black line represents an arbitrarily long path in this example.
Additionally, series-parallel graphs of bounded depth are a generalization of another well known class of structures in physics known as melonic graphs [1, 3]. Those graphs are used in modeling condensed matter and black holes through the Sachdev–Ye–Kitaev Model (SYK) and its generalization the Carrozza, Tanasa, Klebanov, and Tarnopolsky (CTKT) model. Melonic graphs are built recursively using an operation called “bananification” ([1] did not invent it but uses it), this operation creates series-parallel graphs with some 0ptthat corresponds to the order of the melonic graph.
On the other hand, the depth parameter that we consider captures exactly the depth of the tree decomposition of the given series-parallel graph. This draws a parallel to the well-studied structural parameter tree-depth [33], which can be equivalently seen to capture the width of the tree decomposition together with its depth. As series-parallel graphs are of treewidth 2, series-parallel graphs of bounded depth are thus special cases of graphs of bounded tree-depth.
In 2006, Emek and Peleg [15] presented an algorithm that probabilistically embeds series-parallel graphs into distributions over their spanning trees with expected distortion $\mathcal{O}(\log(n))$. Their approach constructs a spanning tree recursively from the composition trace of the graph, ensuring that, in expectation, distances between any pair of vertices remain within a logarithmic factor of the original graph.
We modify the analysis by focusing on the 0ptof the series-parallel graph, in other words, it is the depth of its composition tree as defined in [15]. Formally, this is the number of times the compositions in the trace alternate between Series and Parallel. Note that the 0ptof a series-parallel graph can grow linearly in the worst case (it is upper bounded by $\frac{n}{2}$. However, our claim is not to have a better approximation for all series-parallel graphs but simply that for any fixed 0pt, all series-parallel graph of that 0pthave an efficient approximation. As we discussed in the introduction, even series-parallel graphs of depth $\mathcal{O}(\log n)$ pose a serious obstacle towards better approximation for Requirement Cut.
[…]
Before giving the proof and our statement, it is good to give some context and explanations regarding Emek and Peleg’s algorithm. Their algorithm is called Construct_Tree. Construct_Tree takes as input a series-parallel graph $G$ with two designated terminal vertices $x$ and $y$. The algorithm returns a spanning tree $T$ of $G$. The goal is to ensure that for any pair of vertices, the expected stretch is small between the shortest distance in $G$ and the distance in $T$ is small.
The algorithm operates recursively, using the composition tree (or decomposition trace) of $G$, which reflects how the graph was built from smaller series-parallel components via series and parallel composition. It handles these two cases differently:
  • Series composition: If $G$ is obtained by gluing smaller graphs $G_{1},\ldots,G_{k}$ in series, the algorithm recursively constructs a spanning tree $T_{j}$ for each $G_{j}$, then glues them together in series to form $T$. This process does not increase stretch for any pair $(u,v)$ contained in any $G_{i}$.
  • Parallel composition: If $G$ offers multiple subgraphs $G_{1},\ldots,G_{k}$ that each connect $x$ to $y$, the algorithm first builds a spanning tree $T_{j}$ for each $G_{j}$. It identifies the shortest $(x,y)$-path $P_{k^{*}}$ among them, and keeps $T_{k^{*}}$ untouched. For the other $T_{j}$ (with $j\neq k$), it removes one edge chosen uniformly at random from the path $P_{j}$ connecting $x$ and $y$ in $T_{j}$. This breaks the cycles and ensures the union of all modified trees still forms a spanning tree of $G$. This step is the one that might increase the stretch for some $(u,v)$ contained in a $G_{i}$.
The number of compositions is bounded by $m$ since it is the 0ptof the composition trace that we use to compose our graph. To measure the stretch, Emek and Peleg introduce several variables: $\{a_{i}\}_{i=1}^{m}$, $\{b_{i}\}_{i=1}^{m}$ and $d$ positive integers satisfying:
  • $b_{1}=1$,
  • $b_{i+1}\leq a_{i}+b_{i}$ for all $1\leq i<m$,
  • $c_{i}=\sum_{j=1}^{i-1}a_{i}$,
  • $d\leq a_{m}+b_{m}$,
  • $\sum_{i=1}^{m}a_{i}\leq n$.
Intuitively those variables each mean something different in the construction with respect to a pair of vertices $u,v\in G_{j}$, a graph that does not contain $P_{k}$ the shortest $(x,y)$-path. During a parallel composition operation, $d(u,v)$ might increase if an edge in the shortest $(u,v)$-path is removed:
  • $b_{i}$ is the length of the overlap between the shortest $(u,v)$-path and the $(x,y)$-path in $G_{j}$, $P_{j}$. This is the size of the "risk" portion of $P_{j}$.
  • $a_{i}$ is the length of the rest of $P_{j}$.
  • $c_{i}$ is $d(u,v)-b_{i}$, it is what will remain of the $d(u,v)$ no matter what.
  • $d$ is simply $d(x,y)$.
We will now show how to bound the expected distorsion of Emek and Peleg’s algorithm by an $\mathcal{O}(m)$ factor instead of a $\mathcal{O}(\log(n))$.
Theorem 5.1.
Let $\mathcal{T}$ be the output of Algorithm Construct_Tree [15] when invoked on an series-parallel graph G. Then $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(m),\forall(u,v)\in E(G)$. Where $m$ is the depth of the composition trace of $G$.
Then, we have for $\mathds{E}[dist_{\mathcal{T}}(u,v)]$, the distorsion ([15] Lemma 4.1):
$\mathds{E}[dist_{\mathcal{T}}(u,v)]=\sum\limits_{i=1}^{m}\frac{a_{i}}{a_{i}+b_{i}}\left(\prod\limits_{j=1}^{i-1}\frac{b_{j}}{a_{j}+b_{j}}\right)(b_{i}+c_{i})+\left(\prod\limits_{i=1}^{m}\frac{b_{i}}{a_{i}+b_{i}}\right)(a_{m}+c_{m}+d_{m})$
In their paper [15], Emek and Peleg show that $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(\log n)$ (Lemma 4.2), meaning that the expected stretch of each edge is at most logarithmic in $n$. That is:
$\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(\log n)\quad\text{for all }(u,v)\in E(G)$
We adopt the same notations as them and with a different analysis obtain a different result depending on $m$: $\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\mathcal{O}(m),\forall(u,v)\in E(G)$
$\mathds{E}[dist_{\mathcal{T}}(u,v)]=\sum\limits_{i=1}^{m}\frac{a_{i}}{a_{i}+b_{i}}\left(\prod\limits_{j=1}^{i-1}\frac{b_{j}}{a_{j}+b_{j}}\right)(b_{i}+c_{i})+\left(\prod\limits_{i=1}^{m}\frac{b_{i}}{a_{i}+b_{i}}\right)(a_{m}+c_{m}+d_{m})$
We first rewrite it using $b_{i+1}\leq bi+ai$ which implies that $b_{i}\leq 1+\sum_{j=1}^{i-1}a_{i}$
$\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq\frac{a_{1}}{a_{1}+1}+\sum\limits_{i=2}^{m}\frac{a_{i}}{1+\sum_{j=1}^{i}a_{j}}\left(\frac{1}{1+\sum_{j=1}^{i-1}a_{j}}\right)(1+\sum\limits_{j=1}^{i-1}a_{j}+c_{i})+\left(\frac{(a_{m}+c_{m}+d_{m})}{1+\sum_{i=1}^{m}a_{i}}\right)$
We replace $c_{i}$, using $c_{i}=a_{1}+...+a_{i-1}$
$=\frac{a_{1}}{a_{1}+1}+\sum\limits_{i=2}^{m}\frac{a_{i}}{1+\sum_{j=1}^{i}a_{j}}\left(\frac{1+2\cdot\sum_{j=1}^{i-1}a_{j}}{1+\sum_{j=1}^{i-1}a_{j}}\right)+\left(\frac{d_{m}+\sum_{i=1}^{m}a_{i}}{1+\sum_{i=1}^{m}a_{i}}\right)$
And now we upper bound $d_{m}$ by using $d_{m}\leq 1+\sum_{i=1}^{m}a_{i}$
$\leq\frac{a_{1}}{a_{1}+1}+\sum\limits_{i=2}^{m}\frac{a_{i}}{1+\sum_{j=1}^{i}a_{j}}\left(\frac{1+2\cdot\sum_{j=1}^{i-1}a_{j}}{1+\sum_{j=1}^{i-1}a_{j}}\right)+\left(\frac{1+2\cdot\sum_{i=1}^{m}a_{i}}{1+\sum_{i=1}^{m}a_{i}}\right)$
The first sum has m terms each smaller than 1 and each of them are multiplied by a quantity smaller than 2. The right most term is at most 2.
$\mathds{E}[dist_{\mathcal{T}}(u,v)]\leq 2m+2$
This series of calculations show that for a given 0pt, all series-parallel graphs of that depth can be embedded into a distribution of spanning trees that give a constant expected distorsion for any pair of vertices $(u,v)$. This fact is interesting on its own theoretically and in our case, we can use it to get an $\mathcal{O}(\log(g))$-approximation for the Requirement Cut problem on series-parallel graphs of bounded 0pt.
6 Conclusion
The Requirement Cut problem generalizes numerous classical cut problems, and has resisted single-logarithmic approximations on general graphs due to unavoidable lower bounds like the cost of tree embeddings. In this work, we identify two structural parameters—the number of (minimal) Steiner trees and the depth of series-parallel graphs—that allow us to break the “double-log” barrier.
Our main takeaway is that graphs with a polynomial number of spanning trees, or more generally instances with a polynomial number of minimal Steiner trees, admit efficient LP-rounding algorithms with approximation ratio $\mathcal{O}(\log(n))$. Additionally series-parallel graphs of bounded depth can be approximated in $\mathcal{O}(\log(g))$, extending known results for trees to a broader class. These findings broaden the landscape of graphs for which single-log approximations are possible which was previously limited to trees.
We believe our work represents a natural limit of improvement (up to constant factors) for $LP$ rounding techniques based on randomized thresholding. Future progress may require fundamentally different approaches. Nonetheless, our results suggest that graphs with polynomially many spanning trees and bounded-depth constructions behave in a “tree-like” manner, opening the door to improved approximation algorithms for other connectivity problems in such graphs.
Appendix A Appendix
A.1 Chernoff Bound Details
We take the Chernoff bound presented in Lemma 4.6 and detail each step here:
$\ \mathbb{P}\left(Y_{t}<(1-\delta)\mathds{E}[Y_{t}]\right)\leq e^{-\frac{\delta^{2}}{2}\mathds{E}[Y_{t}]},\quad\text{for }0<\delta<1.$
Substituting $\delta=1-\frac{r^{\prime}-2}{\mathds{E}[Y_{t}]}$, we derive:
$\mathbb{P}(Y_{t}\leq r^{\prime}-2)\leq e^{-\frac{(1-\frac{r^{\prime}-2}{\mathds{E}[Y_{t}]})^{2}}{2}\mathds{E}[Y_{t}]}.$
Expanding the exponent:
$1-\frac{r^{\prime}-2}{\mathds{E}[Y_{t}]}=\frac{\mathds{E}[Y_{t}]-r^{\prime}-2}{\mathds{E}[Y_{t}]},$
$(1-\frac{r^{\prime}-2}{\mathds{E}[Y_{t}]})^{2}=\frac{(\mathds{E}[Y_{t}]-r^{\prime}-2)^{2}}{\mathds{E}[Y_{t}]^{2}}.$
Substituting into the exponent:
$\exp({-\frac{(\mathds{E}[Y_{t}]-r^{\prime}-2)^{2}}{2\mathds{E}[Y_{t}]}})$
Using the lower bound on $\mathds{E}[Y_{t}]$:
$\mathds{E}[Y_{t}]\geq\frac{r^{\prime}-1}{\alpha}$
Using this bound to upper bound the original expression:
Note: This upper-bounding is possible because the function $f(x)=\exp\left(-\frac{(x-r^{\prime}-2)^{2}}{2x}\right)$ is decreasing for any $x\geq r^{\prime}-2$ which is the case for our case.
$\exp\left(-\frac{(\mathds{E}[Y_{t}]-r^{\prime}-2)^{2}}{2\mathds{E}[Y_{t}]}\right)\leq\exp\left(-\frac{\left(\frac{r^{\prime}-1}{\alpha}-r^{\prime}-2\right)^{2}}{2\cdot\frac{r^{\prime}-1}{\alpha}}\right)$
Simplifying the numerator:
$\frac{r^{\prime}-1}{\alpha}-r^{\prime}-2=\frac{r^{\prime}-1-\alpha(r^{\prime}+2)}{\alpha}$
Squaring the numerator:
$\left(\frac{r^{\prime}-1-\alpha(r^{\prime}+2)}{\alpha}\right)^{2}=\frac{(r^{\prime}-1-\alpha(r^{\prime}+2))^{2}}{\alpha^{2}}$
Putting this into the full expression:
$\exp\left(-\frac{(r^{\prime}-1-\alpha(r^{\prime}+2))^{2}}{2\cdot\frac{r^{\prime}-1}{\alpha}\cdot\alpha^{2}}\right)=\exp\left(-\frac{(r^{\prime}-1-\alpha(r^{\prime}+2))^{2}}{2\alpha(r^{\prime}-1)}\right)$
Finally, we know that $r^{\prime}\geq 2$ and we want to substitute. However, we have to make sure that this correctly bounds our term. Similarly to before this upper bounding is indeed possible because $f_{\alpha}(x)=\exp\left(-\frac{(x-1-\alpha(x+2))^{2}}{2\alpha(x-1)}\right)$ admits only one critical points for $x\geq 2$ and it is $x^{*}=\frac{1+2\alpha}{1-\alpha}$. The function increases before that and decreases after. We want it to be decreasing for $x\geq 2$ and so we have to restrict $\alpha$ to $0<\alpha\leq 0.25$ and $x\geq 2$. This is no problem since we can decide on $c$ and make $\alpha$ small enough.
$\exp\left(-\frac{(r^{\prime}-1-\alpha(r^{\prime}+2))^{2}}{2\alpha(r^{\prime}-1)}\right)\leq\exp\left(-\frac{(1-4\alpha)^{2}}{2\alpha}\right)$
$\leq\exp\left(-\frac{16\alpha^{2}-8\alpha+1}{2\alpha}\right)$
Substitute $\alpha$ for $\frac{1}{c\cdot\log(\sigma)}$:
$\exp\left(-\frac{\frac{16}{(c\cdot\log(\sigma))^{2}}-\frac{8}{c\cdot\log(\sigma)}+1}{\frac{2}{c\cdot\log(\sigma)}}\right)=\exp\left(-\frac{\frac{16}{(c\cdot\log(\sigma))}-8+c\cdot\log(\sigma)}{2}\right)$
Here we simply distribute the division by two and then remove a negative term.
$=\exp\left(4-\frac{8}{c\cdot\log(\sigma)}-\frac{c\cdot\log(\sigma)}{2}\right)\leq\exp\left(4-\frac{c\cdot\log(\sigma)}{2}\right)$
$=\frac{\exp(4)}{\sigma^{\frac{c}{2}}}$
This achieves the calculations. You can join back the proof and discussion on $\alpha$ (Section 4.2).